OpenStudy (anonymous):

Use a double-angle identity to find the exact value of the trigonometric expression. cos 420°

OpenStudy (mathmate):

You can always use a double-angle identity for something. Here it is simple enough to use cos(420)=cos(420-360)=cos(60)=1/2

OpenStudy (anonymous):

one to go

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@EmilyJernigan i told you wrong last time it should have been positive, not negative sorry if you can change it, do

OpenStudy (anonymous):

Its fine. Thats what i thought!